OpenStudy (anonymous):

integral of (x+1)/(x^2-5x+6)dx

OpenStudy (anonymous):

OK So What problem are you Facing in this integral ?....

OpenStudy (anonymous):

I don't know how to approach it.

OpenStudy (anonymous):

how do I begin solving it?

OpenStudy (anonymous):

In this Integral... Try to write the above Numerator eq into Denominator eq ...

OpenStudy (anonymous):

x + 1 is Linear.... but x^2 - 5x +6 is Quadratic ! So What if some Kinda of Diiferentiation might work ? Think about it ... Answer lies in Proper Substitution

OpenStudy (anonymous):

yeah, i've thought about this for over an hour now. I really don't know.

OpenStudy (anonymous):

Ok, d( quadratic) A + B = Linear -- dx Here A and B are Constants

OpenStudy (anonymous):

Now Substitute Linear By this Value.... Solve the above Equation to Get A and B

OpenStudy (anonymous):

A d ---(x^2 -5x +6 ) + B = x + 1 dx 2Ax - 5A + B = x + 1 Comparing Co-efficient s 2A = 1 A = 1 -- 2 5A + B =1 5/2 + B =1 B = -3/2 Now REPLACE the Value In the Integral

OpenStudy (anonymous):

I'm Writing the d/dx(quadratic) as D.... A*D + B A*D + B ------- dx = ------- dx -------dx Quadratic Quad Quad Now Put Quad = T d/dx(Quad) =dT/dx Do the substitution .... For dx And Solve :) .... I think this much Help is enough...